# Evaluating the Performance and Robustness of LLMs in Materials Science Q&A and Property Predictions

### v1

@@ Methods
Using MSE-MCQs, we benchmarked a range of both commercial and open-source LLMs, including OpenAI’s gpt-3.5-turbo-0613 [32] and gpt-4-0613 [33], alongside Meta AI’s Llama variants - llama2-70b, llama2-13b, and llama2-7b [34]. The suffixes in the Llama model names denote the number of model parameters where, e.g., llama-70b contains 70 billion parameters. This study primarily utilized fixed-version and open-source models to enhance the reproducibility of the findings.
@@
The prompting strategies used in this evaluation are zero-shot chain-of-thought (CoT) and expert prompting. Zero-shot CoT prompting uses an instructional prompt to enable the model to "think aloud" and step through a reasoning process that leads to an answer, potentially boosting the accuracy in problem-solving tasks [35]. Expert prompting, i.e. instructing an LLM to act as a domain expert, has been shown to influence the model’s responses to be more aligned with expert knowledge and reasoning [36]. The specific implementations of each for their respective tasks are defined below.
@@
Performance Evaluation of Materials Property Prediction
For this study, we used the matbench_steels dataset to evaluate the predictive capabilities of pretrained LLMs, using their ICL capability with few-shot examples. We benchmarked llama2-7b, llama2-13b, mistral-7b [37], gpt-3.5-turbo-0613, and gpt-4-0613 on their abilities to predict the yield strengths given the steel compositions. For performance comparison, a KNN and an RFR model were also implemented. Few-shot learning involves providing the LLM with a few examples of the task at hand, enabling it to learn the pattern and apply it to unseen questions or problems [38]. To use LLMs as predictive models, we fed the “training data” as few-shot examples to the prompt windows of the LLMs. An example of the “training data” prompt is shown below.
@@
composition: Fe0.779 Mn0.000101 Si0.000198 Cr0.114 Ni0.0973 Mo0.000116 Nb0.0000597 Co0.0000941 Al0.00432 Ti0.00359, yield strength: <1125.0>
composition: Fe0.673 Mn0.0000992 Si0.0361 Cr0.167 Ni0.0372 Mo0.000114 Nb0.0000587 Co0.0851 Al0.000404 Ti0.000569, yield strength: <1143.7>
@@
Robustness Analysis of Materials Science Q&A
To evaluate the robustness of LLMs for materials science Q&A, we continued using the MSE-MCQs dataset and tested the gpt-3.5-turbo-0613 model to observe how textual perturbations impacted performance. In this study, no prompting strategy was implemented, and the questions were directly used as the user prompts. We identified different types of “noise” that can be introduced to the MCQs to evaluate the robustness of LLMs. As shown in Table 2, the textual inputs were modified systematically in five different ways, i.e. 1) unit mixing, 2) sentence reordering, 3) synonym replacement, 4) distractive info, and 5) superfluous info.
@@
Superfluous Info Adding materials-science-related superfluous information containing numerical values to the questions To challenge LLMs’ ability to identify relevant data without being misled by additional numeric details that are potentially incorrect or conflicting
Table 2: Types and descriptions of textual degradation applied to LLMs.
@@
Robustness Analysis of Materials Property Prediction
@@
The material descriptions underwent systematic modifications mirroring those applied in the Q&A evaluations, except for unit mixing and synonym replacement. During data preprocessing for LLM-Prop, all numerical values and units, such as bonding distances and angles, are replaced with a [NUM] token, to emphasize the model’s focus on text-based understanding [14]. Unit mixing might disrupt the preprocessing algorithm, and thus was excluded from the analysis. Synonym replacement was excluded because the original terminology was already highly specific and lacked equivalent synonyms. Furthermore, we conducted an ablation study of textual degradation to examine the model’s resilience against structural and length variations in the input data, as well as to explore which aspects of the descriptions the model relies on for predictions. We manipulated the order and fraction of sentences included, testing configurations including 1) original order, which prioritizes the initial information in a description, 2) reverse order, which prioritizes the sentences from the end of a description, 3) random order, shuffling the information, and 4) sides-to-middle, which deprioritized central information. The impact of these textual degradations was quantitatively assessed by measuring the resultant prediction error in MAE.
Data availability
The data required and generated by our code are publicly available at url (to be inserted upon acceptance of the paper).
Code availability
The code for analysis and figure generations in this work is available on GitHub at url (to be inserted upon acceptance of the paper).
Acknowledgements

### v2

@@ 2.1   Performance Evaluation of Materials Science Q&A
Using MSE-MCQs, we benchmarked a range of both commercial and open-source LLMs, including Anthropic’s claude-3.5-sonnet-20240620[28], OpenAI’s gpt-4o-2024-11-20[29], gpt-4-0613 [30], and gpt-3.5-turbo-0613 [31], alongside Meta AI’s Llama variants - llama3.3-70B-instruct[32], llama2-70b-chat, llama2-13b-chat, and llama2-7b-chat [33]. The suffixes in the Llama model names denote the number of model parameters where, e.g., llama-70b contains 70 billion parameters. A reasoning model, DeepSeek-R1[25] was also evaluated. This study primarily utilized fixed-version and open-source models to enhance the reproducibility of the findings.
@@
The prompting strategies used in this evaluation are zero-shot chain-of-thought (CoT) and expert prompting. Zero-shot CoT prompting uses an instructional prompt to enable the model to "think aloud" and step through a reasoning process that leads to an answer, potentially boosting the accuracy in problem-solving tasks [34]. Expert prompting, i.e., instructing an LLM to act as a domain expert, has been shown to influence the model’s responses to be more aligned with expert knowledge and reasoning [35]. The specific implementations of each for their respective tasks are defined below.
@@
2.2   Performance Evaluation of Materials Property Prediction
For this study, we used the matbench_steels dataset to evaluate the predictive capabilities of pretrained LLMs, using their ICL capability with few-shot examples. We benchmarked claude-3.5-sonnet-20240620, gpt-4o-2024-11-20, gpt-3.5-turbo-1106 llama3.3-70B-instruct, mistral-large-2411[36], and Cohere’s command-r7b-12-2024[37] on their abilities to predict the yield strengths given the steel compositions. For performance comparison, a KNN and an RFR model were also implemented. Few-shot learning involves providing the LLM with a few examples of the task at hand, enabling it to learn the pattern and apply it to unseen questions or problems [38]. To use LLMs as predictive models, we fed the “training data” as few-shot examples to the prompt windows of the LLMs. An example of the “training data” prompt is shown below.
@@ 2.2   Performance Evaluation of Materials Property Prediction
composition: Fe0.682 C0.00925 Mn0.000101 Si0.0101 Cr0.134 Ni0.00899 Mo0.0115 V0.000109 Nb0.000479 Co0.143 Al0.000618, yield strength: <1314.2>
composition: Fe0.792 C0.000470 Mn0.000411 Si0.00201 Cr0.0862 Ni0.0980 Mo0.0181 V0.000111 Nb0.0000607 Co0.0000957 Al0.00167 Ti0.000589, yield strength: <1061.7>
@@
2.3   Robustness Analysis of Materials Science Q&A
To evaluate the robustness of LLMs for materials science Q&A, we continued using the MSE-MCQs dataset and tested gpt-4o-2024-11-20 and gpt-3.5-turbo-0613 to observe how textual perturbations impacted performance. In this study, no prompting strategy was implemented, and the questions were directly used as the user prompts. We identified different types of “noise” that can be introduced to the MCQs to evaluate the robustness of LLMs. As shown in Table 1, the textual inputs were modified systematically in five different ways, i.e., 1) unit mixing, 2) sentence reordering, 3) synonym replacement, 4) distractive info, and 5) superfluous info.
Table 1: Types and descriptions of textual degradation applied to LLMs
@@ 2.3   Robustness Analysis of Materials Science Q&A
Superfluous Info Adding materials-science-related superfluous information containing numerical values to the questions To challenge LLMs’ ability to identify relevant data without being misled by additional numeric details
@@
2.4   Robustness Analysis of Materials Property Prediction
@@
The material descriptions underwent systematic modifications mirroring those applied in the Q&A evaluations, except for unit mixing and synonym replacement. Note that, because of the highly templated nature of crystal structure descriptions, superfluous information in this context is better characterized as misleading information rather than simply extraneous text. During data preprocessing for LLM-Prop, all numerical values and units, such as bonding distances and angles, are replaced with a [NUM] token, to emphasize the model’s focus on text-based understanding [15]. Unit mixing might disrupt the preprocessing algorithm, and thus was excluded from the analysis. Synonym replacement was excluded because the original terminology was already highly specific and lacked equivalent synonyms. Furthermore, we conducted a truncation study of textual degradation to examine the model’s resilience against structural and length variations in the input data, as well as to explore which aspects of the descriptions the model relies on for predictions. We manipulated the order and fraction of sentences included, testing configurations including 1) original order, which prioritizes the initial information in a description, 2) reverse order, which prioritizes the sentences from the end of a description, 3) random order, shuffling the information, and 4) sides-to-middle, which deprioritized central information. The impact of these textual degradations was quantitatively assessed by measuring the resultant prediction error in MAE.
3   Results and Discussion
3.1   Performance Evaluation of Materials Science Q&A
The results of the performance evaluation of LLMs on the MSE-MCQs dataset are shown in Figure 2. For each model and setting, three trials were conducted, and the error bars represent the inherent non-determinism in LLMs even at the lowest temperature settings (typically 0). This non-determinism may affect the reliability and reproducibility of the models [42] and is likely a result of stochastic sampling during text generation [43, 44, 45]. We evaluated several advanced models released after late 2024 (i.e., DeepSeek-R1 (reasoning model), claude-3.5-sonnet-20240620, llama3.3-70B-instruct, and gpt-4o-2024-11-20), along with some older models (i.e., gpt-4-0613, gpt-3.5-turbo-0613, and the llama2 variants). Overall, the newer and larger models significantly outperform older models, highlighting substantial recent advancements in LLM capabilities.
Among the evaluated pre-reasoning models, claude-3.5-sonnet-20240620 achieved the highest accuracy across all difficulty levels. Notably, the reasoning model DeepSeek-R1 demonstrated competitive performance, closely matching claude-3.5-sonnet-20240620 on easy and medium questions, and surpassing all models on hard questions. The hard questions predominantly involve complex, multi-step reasoning or advanced mathematical calculations, tasks that typically present substantial challenges for single-pass pre-reasoning models. This superior performance by DeepSeek-R1’s clearly highlights its inherent strength in tasks demanding deeper analytical and mathematical reasoning compared.
The older llama2 models performed at or slightly below the baseline, equivalent to random guessing, while the newer llama3.3-70B-instruct achieved comparable accuracy to gpt-4o-2024-11-20 on easy and medium questions.
Upon implementing the expert prompt (See Expert Prompt), we observe consistent performance improvement across almost all models and question types. The improvement is more significant on the older models, suggesting that the expert prompt can enhance reasoning abilities with weaker baseline capabilities. However, the expert prompt provides minimal benefit for the newer pre-reasoning models on the easy questions, likely because the extensive reasoning process induced by the expert prompt contributes little to the performance on simple conceptual questions that rely primarily on factual recall. Interestingly, DeepSeek-R1 shows no performance improvement on hard questions upon implementing the expert prompt, suggesting that the reasoning capabilities of DeepSeek-R1 are already effectively saturated by its built-in iterative reasoning mechanism, such that additional explicit prompting does not further augment its performance.
We further investigated why the smaller llama2 models (llama-13b-chat and llama-7b-chat) scored lower than the baseline without the expert prompt. Despite being chat models, they sometimes failed to understand the intent when instructions were not provided. Instead of answering, they often attempted to “complete” the questions. Once the expert prompt was implemented, these smaller models could follow the instructions and attempt to solve the questions, in which case the performance improved to around and sometimes above the baseline scores. However, their overall performance remained weak due to their limited skill levels.
Overall, the observed performance trends align with expectations: more recent and larger models consistently demonstrate enhanced capabilities in domain-specific Q&A tasks compared to their predecessors. Additionally, prompt engineering demonstrated effectiveness as a strategy for enhancing model performance when handling more complex questions, especially for older or smaller models with limited baseline capabilities. On the other hand, the reasoning model, DeepSeek-R1, exhibits inherently superior performance in complex analytical and mathematical reasoning tasks, achieving high accuracy even without specialized prompting.
3.2   Performance Evaluation of Materials Property Prediction
Here, we investigate LLM materials property prediction with ICL, utilizing the matbench_steels dataset to predict the yield strength of steels. To explore how the selection of highly relevant training data can be used to enhance LLM performance in property prediction, we conducted a systematic study using nearest-neighbor-boosted ICL. The method developed here involves deliberately selecting training data points based on their representational proximity, meaning that data points with material compositions most similar to the test sample are used as few-shot examples to enhance the model’s performance. Figure 3 shows the prediction performance when LLMs are tasked with predicting yield strength using a) farthest neighbors, b) random neighbors, and c) nearest neighbors. For performance comparison, a k-nearest neighbors (KNN) model and a random forest regressor (RFR) were also implemented. The RFR is explicitly trained using MAGPIE features[39] from the same data points used in ICL. While there is some correlation between material composition and yield strength, composition alone is not a strong predictor of yield strength. The KNN model serves as an additional baseline to demonstrate that the LLMs are not merely averaging or interpolating values from the selected examples but may be identifying implicit patterns in the data. The results are shown in Figure 3.
Figure 3: Prediction performance of various LLMs under different training neighbor settings. Three panels compare the model performance across neighbor selection methods (left to right): a) Farthest Neighbors, b) Random Neighbors, and c) Nearest Neighbors. The TPOT-Mat performance is indicated by a horizontal grey line for benchmarking.
When trained with farthest neighbors, models generally exhibit high MAE with no clear trend as the number of neighbors increases. Most models perform worse than the KNN and RFR models except for claude-3.5-sonnet-20240620, which slightly outperforms the RFR but still exhibits high MAE. These results suggest that, when provided with distant data points, both LLMs and traditional ML models struggle to make valid predictions. This highlights a key challenge in OOD generalization, as training examples that are too dissimilar to the test sample prevent models from capturing meaningful structure-property relationships, leading to higher prediction errors.
For the random neighbors training set, the LLMs’ performance consistently improves as the dataset size increases. This suggests that randomly composed few-shot examples offers a more balanced and diverse learning environment for these models, allowing models to develop more robust generalization. The claude-3.5-sonnet-20240620 and gpt-4o-2024-11-20 models consistently outperform the RFR model as the data size increases, indicating that their more sophisticated architectures and larger training corpora enhance their ability to analyze and interpret complex data relationships more effectively. On the other hand, the smaller and older LLMs (i.e., cohere-command-r7b-12-2024 and gpt-3.5-turbo-1106) exhibit higher MAE values throughout. The random neighbors setting appears to challenge these models to a greater degree, likely due to their smaller scale and pretraining, which limit their ability to generalize effectively to diverse inputs without fine-tuning or additional data processing. However, while their overall performance remains lower, their MAE decreases more significantly with more few-shot examples, suggesting that these LLMs can benefit from more context.
The nearest neighbors represent the most relevant data points in the compositional space to the test points. As expected, the KNN shows an increase in MAE as the number of neighbors grows, since additional (more distant) neighbors are less similar to the predicted sample. If LLMs relied solely on the provided information without additional internal computation, a similar performance decline would be expected. Contrastingly, as the data size increases, all the LLMs show a consistent decrease in MAE and outperform the KNN model after 5 points. The more advanced LLMs (i.e., claude-3.5-sonnet-20240620 and gpt-4o-2024-11-20) consistently outperform the trained RFR model, suggesting that they can capture more complex relationships in the data rather than solely relying on interpolation from the provided examples. Notably, with 25 nearest neighbors, claude-3.5-sonnet-20240620 achieves an MAE of 80.5, nearly matching the best-performing ML model, TPOT-Mat, which achieves an MAE of 79.9 on the matbench_steels dataset [46]. However, it is important to note that this is not a direct comparison, as TPOT-Mat employs a 5-fold nested cross-validation method on Matbench datasets [27], which utilizes 80% of the data and is likely to result in better performance. The results highlight the potential of LLMs in data-lean materials property prediction tasks without the need for feature engineering, particularly given their general-purpose design and lack of task-specific fine-tuning.
The results suggest that pretrained LLMs may exhibit adaptability to new predictive challenges using ICL, particularly when data availability is limited. While their ability to extract patterns from a small number of examples is promising, their performance remains task-dependent and may not generalize across all types of property predictions. A key insight is that LLMs can be potentially valuable in early-stage research or exploratory studies in materials science, where data may be scarce or costly to obtain. One potential use case is active learning, where LLMs help identify the most informative data points for experimental validation, optimizing the data acquisition process and reducing the number of required experiments while still achieving meaningful insights. However, as the number of data points increases, most LLMs suffer from limited prompt windows, which make such applications computationally expensive or impossible, in which case fine-tuned LLMs and traditional machine learning models with dedicated training may be more effective.
To investigate the model’s predictive behaviors under these different settings, we analyzed the parity plots of the claude-3.5-sonnet-20240620’s predictions when utilizing 25 neighboring data points, as shown in Figure 4. Alongside these plots, the figure also includes histograms of the top five most frequently predicted yield strength values, to investigate whether the model is merely guessing a few commonly present values (shown as the red points in the parity plots). This behavior is known as “mode collapse”, whereby a generative model can favor a certain output due to overfitting to its pretraining data or lack of generalization capability [47]. Understanding mode collapse is crucial for evaluating the robustness of LLMs because it directly impacts the model’s reliability and utility in practical applications. By identifying the mode collapse behavior, one can evaluate the validity of those predictions and potentially improve the performance.
Figure 4: Parity plots (top) and associated histograms (bottom) of highlighted modes of claude-3.5-sonnet-20240620 with 25 neighboring data points under different training neighbor settings: a) Farthest Neighbors, b) Random Neighbors, and c) Nearest Neighbors, from left to right.
In the farthest neighbors setting, the red points in the figure form horizontal lines, indicating that the model frequently predicts the same yield strength values regardless of composition.This suggests that it fails to capture the underlying relationship between composition and yield strength effectively. The histogram further reveals a strong mode collapse behavior, with the model repeatedly predicting a set of values. This suggests that the model may be defaulting to a “safe” prediction range (possibly from the pretraining data) when provided with less relevant examples. This aligns with the shortcut learning behavior observed in LLMs, where models rely on superficial correlations rather than learning meaningful patterns from the data [48]. Instead of extrapolating from compositional trends, the model may be leveraging spurious cues from its training distribution, leading to repetitive and less informative predictions. In the random neighbors setting, the model shows better overall performance and a reduced mode collapse behavior. This suggests that introducing more variability into the training data helps the model to better understand the underlying patterns that predict yield strength. The nearest neighbors setting exhibits the best performance suggesting that higher training data proximity can lead to more accurate predictions. The mode collapse behavior is significantly reduced compared to the farthest neighbors and random neighbors, showing a greater diversity in the model’s output.
The observations show varying degrees of mode collapse based on the proximity of prompt examples to the test point. For instance, when provided with more closely related training data, the model exhibits stronger predictive signals, reinforcing its dependence on contextual cues rather than learned generalization. The results from Figure 3 and 4 suggest that LLMs do not develop an intrinsic understanding of structure-property relationships but instead rely heavily on contextual information from the prompt. While they can be effective in-distribution predictors when guided by well-curated few-shot examples, their utility in extrapolative property prediction remains limited. However, in the context of active learning, the mode collapse behavior could serve as an unsupervised indicator of prediction validity, allowing for a self-diagnostic approach to assess whether the model is making meaningful predictions. Since mode collapse manifests as a loss of predictive diversity, its prevalence across test samples may highlight regions of high model uncertainty where additional experimental validation is most needed.
3.3   Robustness Analysis of Materials Science Q&A
In Figure 5, we present the outcomes of the robustness assessment of gpt-3.5-turbo-0613 and gpt-4o-2024-11-20 when confronted with different types of textual modifications to the MSE-MCQs (see Table 1), to evaluate its stability to various types of "noise". The comparison of these two models further demonstrates how the evolution of LLMs has influenced their robustness, contextual understanding, and overall reliability in processing modified inputs.
Figure 5: Robustness analysis of a) gpt-3.5-turbo-0613 and b) gpt-4o-2024-11-20 in materials science Q&A using MSE-MCQs dataset. The error bars represent the standard deviation due to LLMs’ non-determinism (Original, Synonym Replacement, Distractive Info) and the randomness (Sentence Reordering, Unit Mixing, Superfluous Info) introduced to the questions.
Ranking by the degradation severity on the easy-level questions for gpt-3.5-turbo-0613, sentence reordering has the least performance drop, followed by synonym replacement, distractive info, unit mixing, and superfluous info. The performance on the hard-level questions is close to the baseline score, indicating that gpt-3.5-turbo-0613 struggles with complex queries regardless of textual modifications, and will not be discussed in detail. The larger error bars in medium and hard questions suggest that LLMs tend to generate more varied responses to complex and lengthy queries. In contrast, the newer and more advanced model, gpt-4o-2024-11-20, shows minimal degradation on the easy-level questions, maintaining an accuracy above 0.8, except for unit mixing. This suggests that the model is better at handling text changes and more robust than its predecessor. However, performance degradation becomes more noticeable on medium and hard questions. Sentence reordering, unit mixing, and superfluous info have the most significant impact on gpt-4o-2024-11-20 at these difficulty levels.
Sentence reordering has little effect on the performance of both gpt-3.5-turbo-0613 and gpt-4o-2024-11-20 on easy-level questions, indicating that both models can effectively parse and extract key information even when the natural flow of a question is altered. However, the impact becomes more significant on medium and hard-level questions, where reordering appears to disrupt comprehension more significantly. This suggests that while both models exhibit strong syntactic flexibility in simpler cases, they may rely more heavily on common question structures when dealing with more complex queries.
The slight performance drop with synonym replacement suggests that both gpt-3.5-turbo-0613 and gpt-4o-2024-11-20 are somewhat sensitive to changes in terminology, leading to occasional inconsistencies in their responses. This reveals LLMs’ reliance on specific terminologies for recognition and comprehension in materials science. In contrast to humans, who can grasp the conceptual continuity behind varied expressions for flexible cognition, LLMs’ struggles with synonym replacement emphasize the need for advanced training that prioritizes semantic networks over mere word recognition [49].
Introducing distractive information simulates a real-world scenario where irrelevant data often accompanies critical information, requiring sharp focus and analytical precision. Improving LLMs’ ability to filter out irrelevant information is crucial for more effective information retrieval, problem-solving, and data interpretation [50]. The results indicate that both models experience slight degradation on easy-level questions, where mostly simple conceptual knowledge is tested. The added non-essential details seemed to divert the LLMs’ "attention", leading to fewer correct answers. Interestingly, distractive info seems to have improved performance on gpt-4o-2024-11-20 on medium and hard questions. This may be rationalized by the possibility that the added information inadvertently help the model by reinforcing key concepts or encouraging deeper contextual reasoning, functioning similarly to prompt engineering in guiding the model toward more accurate responses.
Mixing and converting the units tests LLMs’ abilities to perform numerical reasoning and apply mathematical concepts within a linguistic context. The added complexity introduced by unit mixing degraded the performance of both models, indicating challenges in handling numerical transformations embedded in text. Although some state-of-the-art LLMs support multi-modal applications and function calls to perform calculations [51], accurately identifying and converting units within large text can still be critical. Improving this ability could enhance LLMs’ effectiveness in tasks such as information retrieval, data interpretation, and scientific analysis, where precise numerical reasoning is essential.
Superfluous information differs from distractive information in that it is more relevant to the questions themselves. The extent of performance degradation is likely influenced by the type and relevance of the superfluous information provided. The results show that gpt-3.5-turbo-0613 struggles significantly with superfluous information, experiencing the most severe performance degradation among all modifications. This suggests that it has difficulty filtering out non-essential details, leading to confusion or misinterpretation. In contrast, gpt-4o-2024-11-20 remains largely unaffected on easy and medium-level, but experiences moderate degradation on hard questions. This indicates that while gpt-4o-2024-11-20 demonstrates stronger information selection capabilities, its ability to filter out unnecessary details weakens as question complexity increases. For LLMs, distinguishing the necessary information from merely relevant but non-essential details is a more challenging cognitive process, mirroring advanced human problem-solving. It requires an understanding of the problem’s objective, prioritizing information based on the question, and applying only the information that will lead to the correct conclusion. This highlights a potential area for improvement in LLMs, particularly in their ability to assess and prioritize critical information in complex reasoning tasks.
3.4   Robustness Analysis of Materials Property Prediction
Table 2: Mean Absolute Error (MAE) under different conditions
Original Distractive Info Sentence Reordering Misleading Info
[…]
The impact of sentence reordering increased the MAE by 12.9%, suggesting the model’s reliance on the structured descriptions for accurate predictions. From the previous study on MSE-MCQs degradation, the effect of sentence reordering was less significant, indicating that larger general LLMs, which are trained with more various texts, can exhibit better contextual understanding and are less prone to order changes.
The presence of misleading information, particularly an additional sentence from another material’s description, leads to a 39% increase in MAE. This substantial degradation indicates that while the model can filter out irrelevant distractive noise, it struggles considerably when faced with data that is contextually relevant to the specific prediction task. Notably, this impact arises from the addition of just a single misleading sentence, highlighting the model’s vulnerability to subtle contextual inconsistencies that misdirect its predictions.
To further assess the model’s robustness and determine which description elements are essential for prediction accuracy, we conducted a truncation study that involves altering the orders and lengths of the input description. As shown in Figure 6, the description length is expressed as percent sentence inclusion, ranging from 10% to 100% and MAE is used as a measure of prediction accuracy.
Figure 6: Oreder and length effect of LLM-Prop on prediction performance.
When the number of description sentences is incrementally increased, the MAE rapidly decreases and minimized at 100% sentence inclusion. Interestingly, in the random order, reversed order, and sides-to-middle configurations, the initial MAE at 10% sentence inclusion is notably lower than in the original order, with some configurations achieving nearly double the performance of the original setting. This indicates that the initial sentences may not contain the most useful information for prediction. The MAEs begin to converge around 50% sentence inclusion, beyond which differences become statistically insignificant. Notably, in the random order setting, there is virtually no variation in the MAE when incorporating three different sentence shuffles. This suggests that LLM-Prop can effectively extract key information and deliver consistent predictions, despite variations in the sentence order.
By 40% sentence inclusion, the reversed order yields the lowest MAE, indicating that sentences at the end of descriptions contain crucial predictive information. However, by 50% sentence inclusion, the performance of the reversed order begins to align with that of the random order, suggesting that central information in the descriptions may not be as crucial for prediction accuracy. Since random order includes more initial sentences than reversed order, this suggests that the first sentences may contribute less relevant details, particularly at lower inclusion percentages. Based on these insights, we developed the sides-to-middle approach, aiming to prioritize information at the beginning and the end. This approach consistently outperforms other configurations between 40% and 70% sentence inclusion, achieving the lowest MAE in this range. The error continues to decrease and is optimized at full sentence inclusion being only 5.8% higher than the original setting in MAE. This result suggests that while the original order remains optimal, the contextual framing provided by the beginning and end of descriptions is particularly important for model accuracy.
This truncation study showcases that the fine-tuned model can perform effectively even when provided with significantly reduced prompts. We found that diverging from the training setup (i.e., changing the textual order of the prompt) can sometimes result in improved performance at truncated data volumes. This counterintuitive result suggests that highly templated training or fine-tuning data can lead to unexpected effects. Consequently, this implies two key considerations: 1) training templates should be diverse to prevent models from overfitting to unimportant patterns, and 2) when using a fine-tuned model trained on a specific template, it may not always be optimal to match the template during inference. These insights highlight the potential for optimizing training costs while maintaining performance.
Conclusions
The findings of this study offer crucial insights into the behaviors and limitations of LLMs in materials science domain-specific Q&A and materials property prediction. While the robustness analysis indicates that LLMs can manage certain types of noise with resilience, their performance is significantly challenged under more complex and deceptive conditions, such as when superfluous information is introduced. In Q&A tasks, prompting techniques (e.g., expert prompting, zero-shot chain-of-thought prompting) can sometimes improve the model performance in handling more complex queries, not by unlocking new capabilities, but by increasing the probability of following an expected format. In materials property predictions, we find that in-context learning allows pretrained LLMs to achieve relatively high accuracy in low-data settings when provided with few-shot examples with high proximity to the target material. However, the observed mode collapse behavior, where the model generates repeated outputs despite varying inputs, showcases that providing ineffective few-shot examples (i.e., out-of-distribution data points to the target material) can cause the model to default to a memorized response rather than conditioning its output on the provided prompt. This study also highlights that fine-tuned models can exhibit enhanced performance under truncated data conditions when diverging from the training setup, such as altering the textual order. While this suggests an unexpected level of robustness, it also exposes the risks associated with fixed templating during fine-tuning, suggesting that users should be cautious about strictly matching the training templates during inference. This study highlights the challenges and limitations of using LLMs in materials science, emphasizing the importance of better dataset curation, dynamic prompting techniques, and training strategies to enhance LLMs as reliable tools for materials discovery and scientific research.
Author contributions
Conflicts of interest
The authors declare no conflicts of interest.
Data and code availability
The datasets and code for the analyses and figure generations in this work are publicly available on GitHub at url: https://github.com/Toniaac/LLM-MSE-Eval-Robustness.git (DOI: 10.5281/zenodo.15007556). Detailed information on data and code access is provided in the README.md file in the GitHub repository.
Acknowledgments